If the population of a town grew from 40 thousand to 50 thousand in ten years, by what percent did the population grow?
Answer by actuary(112)

The growth in population is 50,000-40,000=10,000.
Express the growth in popupation compared to the starting population, i.e. 10,000/40,000. Reduce this fraction to lowest terms. 10,000/40,000=1/4.
Convert 1/4 to a decimal by dividing 1 by 4. 1/4 = .25.
Convert .25 to percent by multipying by 100 and putting a percent sign after the number.
So 1/4 = .25 = 25%